# The Dynamic, the Static, and the Weak factor models and the analysis of high-dimensional time series

### v1

@@ 6.2 Factors and cross-sectional permutations: exchangeability
If the irrelevance of the cross-sectional ordering is taken into account, thus, rate-weak factors are nowhere—much ado about nothing! Note, moreover, that if $\lambda_{j}^{(n)}$ is bounded as $n\to\infty$, ${\rm E}[\lambda_{j}^{(n)}]={\rm E}[\lambda_{j}^{(j)}]$ is a constant.
@@
7 Undetected factors: “don’t throw out the baby with the bathwater!”
@@
Onatski (2012)’s justification for considering rate-weak factors, however, is more subtle. His claim is that weak-factor asymptotics provide a better approximation in a finite-$(n,T)$ situation where the smallest linearly divergent eigenvalues do not separate well from the bulk of bounded eigenvalues. This is what is to be faced in the empirical exercise shown in Figure 3. He does not require, thus, the presence of a “genuinely rate-weak” factor, but uses rate-weak asymptotics as a tool for improved detection and estimation of these hardly detectable strong factors.
One may wonder, though, whether this is worth the effort. Identifying hardly detectable factors and incorporating them in the common space, indeed, has limited practical impact—provided, however, that factor models are not considered as a dimension-reduction technique throwing away the idiosyncratic—an attitude that is adopted by several authors.
Whether static or dynamic, undetected factors, indeed, are not lost, but wrongly left in the empirical idiosyncratic space. In a dimension-reduction perspective, the common component $\chi_{it}$ is considered an approximation of the high-dimensional observation $X_{it}$ which, being reduced rank, is more tractable than $X_{it}$ itself. The idiosyncratic $\xi_{it}$ then is discarded as if it were a negligible error term—a regrettable terminology used by some authors, mainly in the classical literature. This may be legitimate in specific datasets and is common practice, e.g., in macroeconometrics, where it seems that $\xi_{it}$, indeed, safely can be neglected so that forecasting $X_{i,t+1}$ boils down to forecasting $\chi_{i,t+1}$ (Boivin and Ng, 2006; Luciani, 2014).
In many cases, however, $\xi_{it}$ needs not be small and may be strongly autocorrelated. Its empirical version, moreover, may contain undetected (rate-strong, but weakly influential) factors and discarding it is potentially quite damaging, e.g. in a prediction context. This is often the case in financial applications dealing with panels of volatility measures (Herskovic et al., 2016; Barigozzi and Hallin, 2017, 2020).
Rather than a dimension reduction technique, factor models, actually, should be considered a “divide and conquer” procedure where the common and the idiosyncratic are analyzed via distinct appropriate methods—reduced-rank-process techniques for the common, taking advantage of the strong cross-correlations, and componentwise or sparse techniques for the idiosyncratic, where no advantage can be gained from the negligible potential cross-correlations. The resulting analyses, at the end of the day, are to be brought back together: for instance, once forecasts ${\widehat{\chi}}_{i,t+1}$ and ${\widehat{\xi}}_{i,t+1}$ of the common and the idiosyncratic values at time $t+1$ have been obtained, their sum ${\widehat{\chi}}_{i,t+1}+{\widehat{\xi}}_{i,t+1}\eqqcolon\widehat{X}_{i,t+1}$ should be used as a predictor for $X_{i,t+1}$. This point is clearly made by Fan et al. (2023) in the context of approximate static factor models and by Barigozzi et al. (2024) for the GDFM.

### v3

@@ 7.2 Factors and cross-sectional permutations: exchangeability
If the irrelevance of the cross-sectional ordering is taken into account, thus, rate-weak factors are nowhere—much ado about nothing!
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: FRED-MD dataset (McCracken and Ng, 2016), $r=8$. Evolution, for increasing $n=1,\ldots,n_{0}$, of the eigenvalues $\lambda_{X;k}^{(n)}$, $k=1,\ldots,8$ (all normalized by $\lambda_{X;1}^{(n_{0})}$) under 100 random cross-sectional permutations.
In Figure 3, as a numerical illustration of that fact, we consider the FRED-MD dataset described by McCracken and Ng (2016), which consists of $n_{0}=126$ monthly times series covering both the real and nominal sectors of the U.S. economy and also includes labor market indicators and financial variables (see https://research.stlouisfed.org/econ/mccracken/fred-databases/). The data is transformed to stationarity and missing values are imputed by means of the routines made available by McCracken and Ng (2016), which produces a balanced panel, with a sample running from April 1959 through March 2024—an observation period of $T=780$ time points.
The cross-sectional ordering, in this dataset, is fully arbitrary, hence should be meaningless. Standard methods (Bai and Ng, 2002) for determining the number of static factors point to as much as $\widehat{r}=8$ factors. Figure 3 shows the evolution, for increasing $n=1,\ldots,n_{0}$, of the eigenvalues $\lambda_{X;k}^{(n)}$, $k=1,\ldots,8$ (all normalized by $\lambda_{X;1}^{(n_{0})}$) of ${\bm{\Gamma}}^{(n)}_{X}$ when considering 100 random permutations of the cross-section. The plots show that the evolution of all eigenvalues is well described by a linear (in $n$) divergence, with decreasing (in $k$) slopes. These slopes, for the two or three last eigenvalues, are small and one may be tempted to consider them as rate-weak and speculate about the value $\alpha$ in some sublinear divergence rate $n^{\alpha}$: this is, however, disputable and incompatible with the irrelevance of cross-sectional ordering.
8 Identifying the number of factors
From the above discussions, it is clear that the numbers of factors ($r$ in the static decomposition (5.2)satisfying (5.3) and $q$ in the GDFM decomposition (3.3) satisfying (3.4)) are only asymptotically identified as $n\to\infty$, and coincide with the number of diverging eigenvalues of the covariance and spectral density matrices, respectively. In practice, with a finite-$(n,T)$ realization $\{X_{it}|\,i=1,\ldots,n,\ t=1,\ldots,T\}$ of the process $\{X_{it}|\,i\in\leavevmode\nobreak\ \!\mathbb{N},\,\leavevmode\nobreak\ \!t\in\leavevmode\nobreak\ \!\mathbb{Z}\}$, an estimation of these numbers has to be based on estimators, $\widehat{\bm{\Gamma}}^{(n)}_{X}$ or $\widehat{\bm{\Sigma}}^{(n)}_{X}(\theta)$, of these matrices and the corresponding empirical eigenvalues $\widehat{\lambda}^{(n)}_{X;j}$ or $\widehat{\lambda}^{(n)}_{X;j}(\theta)$, $\theta\in[-\pi,\pi]$.
Whether the objective is the identification/estimation of $r$ or $q$, two main strategies can be found in the literature. The most popular one—call it the eigengap detection strategy—is based on the fact that $r$ and $q$ (unless they are zero) are characterized by an arbitrarily large “eigengap” between unbounded and bounded eigenvalues. More precisely, $r$ is the largest value of $j$ such that ${\lambda}^{(n)}_{X;j}-{\lambda}^{(n)}_{X;j+1}\to\infty$ as $n\to\infty$, and $q$ the largest value of $j$ such that $\int_{-\pi}^{\pi}\left\{{\lambda}^{(n)}_{X;j}(\theta)-{\lambda}^{(n)}_{X;j+1}(\theta)\right\}{\rm d}\theta\to\infty$ as $n\to\infty$. A sensible strategy, thus, consists in selecting $\widehat{r}^{\,(n,T)}$ and $\widehat{q}^{\,(n,T)}$ as the maximal $j$’s yielding a “large eigengap” between the empirical counterparts $\widehat{\lambda}^{(n)}_{X;j}-\widehat{\lambda}^{(n)}_{X;j+1}$ and $\int_{-\pi}^{\pi}\left\{\widehat{\lambda}^{(n)}_{X;j}(\theta)-\widehat{\lambda}^{(n)}_{X;j+1}(\theta)\right\}{\rm d}\theta$ of these quantities (note that, as functions of the frequency $\theta\in[-\pi,\pi]$, these $\widehat{\lambda}^{(n)}_{X;j}(\theta)$ do not cross). Depending on what is meant with a “large eigengap,” several implementations of this strategy have been proposed. The prototype for the identification of $r$ is the information criterion method of Bai and Ng (2002), which is also the most popular, followed by many others among which Onatski (2010), based on the asymptotic distribution of the $\widehat{\lambda}^{(n)}_{X;j}$’s, and the eigenvalue ratio approaches by Ahn and Horenstein (2013) and Trapani (2018)—to quote only a few.
A different strategy is based on the fact that the divergence of eigenvalues is a property of sequences of the type ${\lambda}^{(m)}_{X;j},{\lambda}^{(m+1)}_{X;j},\ldots$ or ${\lambda}^{(m)}_{X;j}(\theta),{\lambda}^{(m+1)}_{X;j}(\theta),\ldots$ This strategy therefore considers the corresponding empirical trajectories rather than restricting to the “terminal” empirical values $\widehat{\lambda}^{(n)}_{X;j}$ and $\widehat{\lambda}^{(n)}_{X;j}(\theta)$ and was first proposed by Hallin and Liška (2007) for the identification of $q$ in the GDFM. That method is designed to minimize, over a grid of increasing values of the cross-sectional dimension, the trace of the idiosyncratic spectral density subject to a penalty which is tuned by means of a positiveconstant $c\in\leavevmode\nobreak\ \![0,c_{\max}]$ where $c_{\max}>0$ is some pre-specified maximal value. More precisey, for a grid ${\mathfrak{G}}^{(n)}\coloneqq\{0=m_{1}<\ldots<m_{\nu}<\ldots<m_{N(n)}=n\}$ of increasing cross-sectional dimensions and a grid ${\mathfrak{C}}\coloneqq\{0<c_{1}<\ldots<c_{\ell}<\ldots c_{L}=c_{\max}\}$ of increasing tuning constants, this method considers, for each $m\in{\mathfrak{G}}^{(n)}$, fixed $n$, and $c\in{\mathfrak{C}}$, the minimization problem
$\min_{1\leq k\leq q_{\max}}\text{IC}^{(m)}(k;c),\;\text{ where }\;\text{IC}^{(m)}(k;c)\coloneqq\sum_{j=k+1}^{m}\int_{-\pi}^{\pi}\widehat{\lambda}^{(m)}_{X;j}(\theta)\mathrm{d}\theta+kc\;\mathrm{p}(m)$
for some penalty function $\mathrm{p}$ and pre-specified $q_{\max}$. Solving these problems yields, for each $c\in{\mathfrak{C}}$, a sequence $\check{q}_{c}^{(n)}(m)$, $m\in\leavevmode\nobreak\ {\mathfrak{G}}^{(n)}$ of possible numbers of factors. For adequate penalty function $\mathrm{p}$, it can be shown that all these sequences, irrespective of $c$, are consistent as $n\to\infty$. The choice of an “optimal penalty” $c$, denote it as $c^{*}$, can be based on the fact that overpenalization ($c$ too large) yields consistency from below while underpenalization yields consistency from above; this suggests selecting $c^{*}$ as any value of $c$ achieving, as $m$ ranges over ${\mathfrak{G}}^{(n)}$, the most “stable” trajectory of $\check{q}_{c}^{(n)}(m)$ values (typically, an interval of $[0,c_{\max}$). The selected number of factors then is $\widehat{q}^{\,(n,T)}\coloneqq\check{q}_{c^{*}}^{(n)}(n)$. Hallin and Liška (2007) establish the consistency consistency under linearly divergent eigenvalues; Onatski (2024) shows that this consistency still holds when the factors are extremely (rate-)weak.
An analogue of this method for the static case and the estimation of $r$ is proposed by Alessi et al. (2010), who refine the information criterion approach of Bai and Ng (2002) by introducing, as in Hallin and Liška (2007), a tuning of the penalty which enhances its finite-sample performance.
The performance of these methods, however, crucially depends on the following two issues:
  1. (a)
    the quality of $\widehat{\lambda}^{(n)}_{X;j}$ and $\widehat{\lambda}^{(n)}_{X;j}(\theta)$ as estimators of ${\lambda}^{(n)}_{X;j}$ and ${\lambda}^{(n)}_{X;j}(\theta)$, and
  2. (b)
    the quality of ${\lambda}^{(n)}_{X;j}$ and ${\lambda}^{(n)}_{X;j}(\theta)$ as approximations of the corresponding eigenvalues ${\lambda}_{X;j}$ and ${\lambda}_{X;j}(\theta)$ of the infinite-dimensional covariance and spectral density matrices ${\bm{\Sigma}}_{X}$ and ${\bm{\Sigma}}_{X}(\theta)$ of the observed process $\{X_{it}|\,i\in\mathbb{N},\ t\in\mathbb{Z}\}$.
The former (a) is an estimation problem, the latter a population feature on which we have no control. As far as (a) is concerned, many results on the estimation of large covariance ands spectral density matrices (and their eigenvalues) are available in the literature, and we will not discuss this here—see, e.g., Wu and Zaffaroni (2018) and Zhang and Wu (2021) for a modern treatment.
A third point, moreover, is to be taken into account in view of the irrelevance of cross-sectional ordering and the arguments developed in Section 7:
  1. (c)
    the selected values $\widehat{q}^{\,(n,T)}$ and $\widehat{r}^{\,(n,T)}$ should be invariant under cross-sectional permutations.
All “eigengap detection” methods naturally satisfy this invariance requirement since $\widehat{\lambda}^{(n)}_{X;j}$ and $\widehat{\lambda}^{(n)}_{X;j}(\theta)$ do for all $j$. But the Hallin and Liška (2007) and Alessi et al. (2010) methods do not because the sequences ${\lambda}^{(m)}_{X;j}(\theta)$, $m=j,\ldots,n$ very much depend on the cross-sectional ordering. A natural way to eliminate this dependence in the Hallin and Liška (2007) method consists in averaging the $\widehat{\lambda}^{(m)}_{X;j}(\theta)$’s over several random cross-sectional permutations. That average typically stabilizes very rapidly with the number of permutations and, when performed on the resulting sequences $\overline{\lambda}^{(m)}_{X;j}(\theta)$, $m=j,\ldots,n$, the Hallin and Liška (2007) method satisfies (c). The same remark holds, and the same averaging is in order for the estimator $\widehat{r}^{\,(n,T)}$ of $r$ based on the Alessi et al. (2010) method and the static eigenvalues $\widehat{\lambda}^{(n)}_{X;j}$.
The literature is mostly focused on the consistency properties, as $(n,T)\to\infty$, of $\widehat{q}^{\,(n,T)}$ as an estimator of $q$ (of $\widehat{r}^{\,(n,T)}$ as an estimator of $r$). However, desirable as it is, consistency is only an asymptotic property, and offers little finite-sample guarantees: both the “eigengap strategy” and the Hallin and Liška (2007) one, for finite $n$ and depending on (b) above, may fail to separate unbounded and bounded eigenvalues: essentially, they will separate the “rapidly diverging” ones (e.g., linearly divergent with “large” slopes) from the “slowly diverging” (linearly divergent with “small” slopes) ones—the latter being wrongly considered as bounded, i.e., idiosyncratic. Despite of consistency, this often leads to finite-$n$ underestimation of the actual value of $q$ and undetected factors that are absorbed into the idiosyncratic components.
A slowly diverging eigenvalue does not imply that the corresponding undetected factor or shock is negligible, though, and this is why, in Section 9, we recommend not to throw idiosyncratic components away, e.g. in forecasting problems.
@@ 8 Identifying the number of factors
9 Undetected factors, idiosyncratic components, orthogonality, and forecasting: “don’t throw out the baby with the bathwater!”
@@
Onatski (2012)’s justification for considering rate-weak factors is more subtle, though. His claim is that weak-factor asymptotics provide a better approximation in a finite-$(n,T)$ situation where the smallest linearly divergent eigenvalues do not separate well from the bulk of bounded eigenvalues. He does not require, thus, the presence of a “genuinely rate-weak” factor, but uses rate-weak asymptotics as a tool for improved detection and estimation of poorly pervasive, hence hardly detectable, strong factors.